The chemical reaction that causes iron to corrode in air is given by in which at 298 K rxn AS-543.7 J/KPART A:

What is the Gibbs free energy for this reaction at 3652 K ? Assume that ?H and ?S do not change with temperature.

PART B:

At what temperature Teq do the forward and reverse corrosion reactions occur in equilibrium?
Answer #1

Part A

use the direct equation

ΔGrxn = ΔH -TΔS

ΔH rxn = -1684 kJ

ΔS rxn = -543.7 J /K convert in to kJ = -0.5437 kJ /K

T = 3652 K

put these values in the above equation

ΔGrxn = -1684 kJ - 3652K (-0.5437 kJ/ K)

ΔGrxn = -1684 kJ + 1985.6 kJ

ΔGrxn = 301.6 kJ

Part B

at equilibrium deltaG = must be zero

ΔGrxn = ΔH -TΔS at equilibrium this reaction will be

ΔH -TΔS = 0

ΔH = TΔS

T = ΔH/ΔS

put this values in this expression

T = - 1684 kJ / -0.5437 kJ/ K

T = 3097.3 K

at this temperature T = 3097.3

lets check this

ΔGrxn = ΔH -TΔS

if you put this temperature value in the above equation you should get the deltaG value zero

ΔGrxn = -1684 kJ -3097.3 K (-0.5437 kJ/K)

ΔGrxn = -1684 kJ +1684 kJ

ΔGrxn = 0